# Bayesian Ego-graph Inference for Networked Multi-Agent Reinforcement Learning

### v3

@@ 1 Introduction
To address this, we propose BayesG, a decentralised graph-based actor–critic framework for networked MARL that learns latent interaction structures via Bayesian inference. As illustrated in Figure 1(b), we begin by introducing a graph-based policy, where each agent conditions its decisions on a stochastic subgraph sampled from a learned distribution over its physical neighbourhood. Each agent operates over an ego-graph, a localized subgraph capturing its immediate neighborhood, enabling context-aware decisions under topological constraints. We formulate this process as Bayesian variational inference of latent graphs, where each agent $i$ infers a binary mask $Z_{i}$ over its local neighbourhood. The mask is treated as a latent variable with posterior $p(Z_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}},D_{i})$, conditioned on the physical subgraph $G^{\text{env}}_{\mathcal{V}_{i}}$ and agent-specific data $D_{i}$ (e.g., neighbour states, trajectories and polices). The posterior is approximated by a variational distribution $q(Z_{i};\phi_{i})$, optimized end-to-end via the evidence lower bound (ELBO).
@@ 3 Preliminaries
We consider networked MARL, over a fixed interaction graph $G=(\mathcal{V},\mathcal{E})$, where each agent $i\in\mathcal{V}$ interacts locally with neighbors $\mathcal{N}_{i}=\{j\mid(i,j)\in\mathcal{E}\}$ and defines a closed neighborhood $\mathcal{V}_{i}:=\mathcal{N}_{i}\cup\{i\}$. At each step, agent $i$ receives a local observation $\tilde{s}_{i}=f(s_{\mathcal{V}_{i}})$, which may include its own state and aggregated information from neighbors, and selects an action $u_{i}\in\mathcal{U}^{i}$ according to a local policy $\pi_{i}(u_{i}\mid\tilde{s}_{i})$. The encoder $f(\cdot)$ captures neighbourhood-level context using neural architectures (e.g., MLPs [7]); specific choices are detailed in Section 5.1.2. We begin with the spatiotemporal MDP formulation, followed by the decentralised actor–critic (A2C) training objective.
@@
$p_{i}(s_{i}^{\prime}\mid s_{\mathcal{V}_{i}},u_{i})=\sum_{u_{\mathcal{N}_{i}}}\left(\prod_{j\in\mathcal{N}_{i}}\pi_{j}(u_{j}\mid f(s_{\mathcal{V}_{j}}))\right)p(s_{i}^{\prime}\mid s_{\mathcal{V}_{i}},u_{i},u_{\mathcal{N}_{i}}),$
@@ 4.1 Graph-based Policy with Latent Interaction Structures
where $G_{\mathcal{V}_{i}}\in\{0,1\}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}$ is a sampled binary adjacency matrix over agent $i$’s closed neighborhood, drawn from a distribution $\rho(\cdot)$ parameterized by $\varphi_{i}$. The graph is restricted to be a subgraph of the environment topology, i.e., $G_{\mathcal{V}_{i}}\in\mathcal{G}^{\text{env}}_{\mathcal{V}_{i}}:=\{G\in\{0,1\}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}\mid G\preceq G^{\mathrm{env}}_{\mathcal{V}_{i}}\},$ ensuring agents can only reason over physically permitted connections. The function $\tilde{f}_{i}(\cdot)$ denotes a graph-conditioned encoder over $s_{\mathcal{V}_{i}}$ and $G_{\mathcal{V}_{i}}$. $\tilde{\pi}_{i}(\cdot\mid\cdot;\theta_{i})$ is the action-selection policy conditioned on the encoded input, parameterized by $\theta_{i}$.
This formulation enables each agent to stochastically select contextually relevant neighbors for coordination, driven by local observations and task-specific demands. In real-world settings such as traffic signal control, this allows agents to dynamically emphasize high-impact neighbors (e.g., congested junctions) and suppress irrelevant ones, improving both communication efficiency and decision quality while remaining consistent with the underlying road network.
@@ 4.2 BayesG: Variational Latent Ego-graph Inference for Policy Optimisation
We formalise the learning of the latent edge mask $Z_{i}$ as a Bayesian inference problem. Recall that the graph distribution $\rho(G_{\mathcal{V}_{i}}\mid s_{\mathcal{V}_{i}})$, introduced in Definition 3, governs which neighbors agent $i$ attends to. We model this via a binary mask $Z_{i}$ applied over the physical neighborhood graph $G^{\text{env}}_{\mathcal{V}_{i}}$, yielding the effective subgraph:
@@
Treating edge masks as latent variables enables agents to capture uncertainty over which neighbors are informative—particularly valuable when noisy or non-stationary dynamics make deterministic graph pruning unreliable.
@@
The variational parameters $\phi_{i}$ are learned by minimizing the Kullback–Leibler (KL) divergence between the approximate posterior $q(Z_{i};\phi_{i})$ and the true Bayesian posterior $p(Z_{i}\mid G^{\mathrm{env}}_{\mathcal{V}_{i}},D_{i})$. This approximation enables each agent to learn a structured distribution over interaction graphs that reflects the most informative subgraphs for policy optimization. Following the standard variational inference framework, this is equivalent (up to an additive constant) to maximizing the evidence lower bound (ELBO), or equivalently, minimizing the negative ELBO:
@@
where the likelihood term $\log p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})$ is instantiated by the graph-conditioned actor loss $\mathcal{L}_{\theta,\varphi}$ defined in Definition 4, which provides task feedback to guide the learning of latent structures.
@@
$\mathcal{L}_{\mathrm{ELBO}}=\mathbb{E}_{q(Z_{i};\phi_{i})}\!\left[-\mathcal{L}_{\theta,\varphi}\!+\!\sum_{j\in\mathcal{N}_{i}}\left((\lambda+\sigma(\phi_{ij}))\log\sigma(\phi_{ij})\!+\!(2\!-\!\lambda-\sigma(\phi_{ij}))\log(1\!-\!\sigma(\phi_{ij}))\right)\right],$ (9)
where $\mathcal{L}_{\theta,\varphi}$ denotes the graph-conditioned policy loss defined in Definition 4, and $\sigma(\phi_{ij})$ is the Bernoulli parameter for each candidate edge. This objective jointly optimizes the policy and the structure of the latent subgraph via variational inference.
@@ 5.1 Experimental Setup
We evaluate BayesG on five benchmark scenarios for adaptive traffic signal control (ATSC), implemented using the SUMO microscopic traffic simulator [38]. Each environment simulates peak-hour traffic, with one MDP step corresponding to a fixed control interval. Agents observe traffic conditions via induction-loop detectors (ILDs), including vehicle density, queue lengths, and waiting times on incoming lanes, and control local traffic signals. The reward is the negative number of halted vehicles, normalized by a fixed scale. These environments naturally satisfy Definition 1: state transitions depend on local and neighbor actions, rewards are localized per agent, and neighborhood structure is fixed by the road network. Detailed MDP mappings are in Appendix B.
@@
NewYork. To assess scalability, we introduce three large-scale scenarios—NewYork33, NewYork51, and NewYork167—comprising 33, 51, and 167 signalized intersections, respectively. These networks are derived from real-world Manhattan layouts (See Appendix C for more details). We use a 20-second control interval (increased to 40 seconds for NewYork167) and simulate 500 steps per episode. Intersections feature heterogeneous phase designs and ILD configurations. States include normalized lane-level metrics and neighborhood-aware features. These environments pose significant challenges for coordination and generalization in decentralized MARL.
@@ 5.1.1 Baselines
ConseNet [16]: A consensus-based method where critics synchronise neighbours’ parameters via local averaging to promote stability and coordination, without exchanging message content.
@@ 5.4 Ablation Studies
Mask Input Features. We further examine how different information types affect the learned mask. Specifically, we use:(1) State: Mask is generated based on neighbor observations $s_{\mathcal{V}_{i}}$. (2) Trajectory: Uses LSTM hidden states $h_{\mathcal{V}_{i}}$, capturing temporal behavior. (3) Policy: Based on neighbor action distributions $\pi_{\mathcal{N}_{i}}$. Results in Figure 5 (right) indicate tha trajectory-based features lead to stronger performance tha state-only, as the incorporate richer temporal context. Policy-based masks also provide meaningful signals, reflecting agents’ behavioral intentions. The best results are achieved when the learned mask uses a combination of these features, as implemented in BayesG.
@@ 6.1 Comparison with Sampling-based Networked MARL
Recent work has explored sampling strategies for scalability in networked MARL, but differs fundamentally in what is sampled, how distributions are determined, and where adaptivity occurs. (1)Sampling targets and distributions.Qu et al. [26] establish convergence results for networked systems with local dependencies on fixed graphs. Lin et al. [23] sample active links from a predefined distribution $\mathcal{D}$ for $\mu$-decay analysis. Anand and Qu [27] use uniform agent subsampling ($k$ out of $n$) for a global controller with mean-field bounds. In contrast, our method learns context-dependent, per-edge distributions: each agent maintains a learned posterior $q(Z_{i};\phi_{i})$ over its 1-hop edges, optimized jointly with the policy via an ELBO. (2) Locus of adaptivity. Lin et al. [23] achieve dynamics-level stochasticity through time-varying link sets with multi-hop influence across time. Anand and Qu [27] vary agent subsets for a global controller. Our approach operates at the policy/representation level: each agent samples a task-adaptive 1-hop subgraph for local decision-making. (3) First-hop inclusion probabilities. Lin et al. [23] use the marginal of $\mathcal{D}$ (predefined, not state-adaptive). Anand and Qu [27] use uniform inclusion ($k/n$) with no edge-specific notion. Our method learns edge-specific probabilities $\Pr[z_{ij}=1]=\sigma(\phi_{ij}(\cdot))$ that are state and trajectory-dependent. These approaches provide valuable convergence guarantees under structural assumptions. Our contribution is complementary: learning which neighbors to attend to, optimized end-to-end with the policy. Extending their theoretical tools to learned edge masks is a promising future direction. A detailed side-by-side comparison is provided in Table 2 (Appendix H).
@@ 6.2 Distinction from Dec-POMDP-based Coordination Graphs
In contrast, Dec-POMDP-based methods such as CASEC [15], DCG [33], DICG [29], SOP-CG [34], and GACG [22] assume global state $s$, global reward $R(s,\mathbf{u})$, joint transitions $p(s^{\prime}|s,\mathbf{u})$, and centralized training with decentralized execution (CTDE). They allow unrestricted graph rewiring unconstrained by physical topology, making them incompatible with physically grounded networked systems.
Our baselines—CommNet [40], NeurComm [7], and LToS [23]—communicate only with physical neighbors, respect fixed topology, and perform fully decentralized learning. While Dec-POMDP coordination graphs are important contributions, we exclude them due to incompatibility with Spatiotemporal-MDP’s decentralized, physically constrained nature.
@@ 7 Conclusion
We presented BayesG, a variational framework for learning latent interaction graphs in networked multi-agent reinforcement learning. By operating over ego-graphs—localized subgraphs centered on each agent—BayesG enables decentralized agents to infer task-adaptive communication structures using local observations and Bayesian inference.
Our approach integrates graph inference with actor–critic training via an ELBO objective, jointly optimizing both the interaction topology and policy. Experiments on five adaptive traffic signal control benchmarks demonstrate that BayesG consistently outperforms strong baselines, achieving superior coordination and reduced congestion. Case studies and ablation results further highlight the efficiency and interpretability of the learned graphs. BayesG offers a principled approach to structure-aware cooperation in decentralized settings and opens up new opportunities for scalable learning in networked systems.
@@ A.3 Derivation of Variational Objective and ELBO for BayesG
$=\mathbb{E}_{q(Z_{i})}\left[\log q(Z_{i})-\log p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})-\log p(Z_{i})\right]+\log p(D_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}}).$ (24)
Since the marginal likelihood $\log p(D_{i}\mid G^{\mathrm{env}}_{\mathcal{V}_{i}})$ is constant with respect to the variational parameters $\phi_{i}$, it can be omitted during optimization. Therefore, minimizing the KL divergence is equivalent (up to an additive constant) to maximizing the following evidence lower bound (ELBO):

### v4

@@ 1 Introduction
To address this, we propose BayesG, a decentralized graph-based actor–critic framework for networked MARL that learns latent interaction structures via Bayesian inference. As illustrated in Figure 1(b), we begin by introducing a graph-based policy, where each agent conditions its decisions on a stochastic subgraph sampled from a learned distribution over its physical neighbourhood. Each agent operates over an ego-graph, a localized subgraph capturing its immediate neighborhood, enabling context-aware decisions under topological constraints. We formulate this process as Bayesian variational inference of latent graphs, where each agent $i$ infers a binary mask $Z_{i}$ over its local neighbourhood. The mask is treated as a latent variable with posterior $p(Z_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}},D_{i})$, conditioned on the physical subgraph $G^{\text{env}}_{\mathcal{V}_{i}}$ and agent-specific data $D_{i}$ (e.g., neighbour states, trajectories and polices). The posterior is approximated by a variational distribution $q(Z_{i};\phi_{i})$, optimized end-to-end via the evidence lower bound (ELBO).
@@ 3 Preliminaries
We consider networked MARL, over a fixed interaction graph $G=(\mathcal{V},\mathcal{E})$, where each agent $i\in\mathcal{V}$ interacts locally with neighbors $\mathcal{N}_{i}=\{j\mid(i,j)\in\mathcal{E}\}$ and defines a closed neighborhood $\mathcal{V}_{i}:=\mathcal{N}_{i}\cup\{i\}$. At each step, agent $i$ receives a local observation $\tilde{s}_{i}=f(s_{\mathcal{V}_{i}})$, which may include its own state and aggregated information from neighbors, and selects an action $u_{i}\in\mathcal{U}^{i}$ according to a local policy $\pi_{i}(u_{i}\mid\tilde{s}_{i})$. The encoder $f(\cdot)$ captures neighbourhood-level context using neural architectures (e.g., MLPs [7]); specific choices are detailed in Section 5.1.2. We begin with the spatiotemporal MDP formulation, followed by the decentralized actor–critic (A2C) training objective.
@@
$p_{i}(s_{i}^{\prime}\mid s_{\mathcal{V}_{i}},u_{i})=\sum_{u_{\mathcal{N}_{i}}}\left(\prod_{j\in\mathcal{N}_{i}}\pi_{j}(u_{j}\mid f(s_{\mathcal{V}_{j}}))\right)p(s_{i}^{\prime}\mid s_{\mathcal{V}_{i}},u_{i},u_{\mathcal{N}_{i}}),$ (1)
@@ 4.1 Graph-based Policy with Latent Interaction Structures
where $G_{\mathcal{V}_{i}}\in\{0,1\}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}$ is a sampled binary adjacency matrix drawn from distribution $\rho(\cdot;\varphi_{i})$ over agent $i$’s closed neighborhood. The graph is restricted to the environment topology: $G_{\mathcal{V}_{i}}\in\mathcal{G}^{\text{env}}_{\mathcal{V}_{i}}:=\{G\in\{0,1\}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}\mid G\preceq G^{\mathrm{env}}_{\mathcal{V}_{i}}\}$, ensuring only physically permitted connections. The function $\tilde{f}_{i}(\cdot)$ is a graph-conditioned encoder, and $\tilde{\pi}_{i}(\cdot\mid\cdot;\theta_{i})$ is the action-selection policy.
This formulation enables agents to stochastically select contextually relevant neighbors, driven by local observations. By formulating coordination as a distribution $\rho(\cdot)$ over subgraphs, this approach provides a general framework for inferring context-dependent interaction structures—allowing agents to explore diverse coordination patterns and adapt to non-stationary dynamics. In traffic signal control, this allows dynamically emphasizing congested junctions while suppressing irrelevant neighbors.
@@ 4.2 BayesG: Variational Latent Ego-graph Inference for Policy Optimisation
Recall that the graph distribution $\rho(G_{\mathcal{V}_{i}}\mid s_{\mathcal{V}_{i}})$, introduced in Definition 3, governs which neighbors agent $i$ attends to. To infer this stochastic distribution over the fixed physical topology, we model it via a binary mask $Z_{i}\in\{0,1\}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}$ applied over the physical neighborhood graph $G^{\text{env}}_{\mathcal{V}_{i}}$, where each entry $z_{ij}=1$ indicates that agent $i$ communicates with neighbor $j$. This yields the effective subgraph:
@@
We formalise the learning of the latent edge mask $Z_{i}$ as a Bayesian inference problem. This is a natural choice as Bayes’ theorem offers a principled way to infer a latent distribution conditioned on the directly observable physical graph $G^{\text{env}}_{\mathcal{V}_{i}}$ and agent-local data $D_{i}$. Treating edge masks as latent variables enables agents to capture uncertainty over coordination structures—particularly valuable when noisy or non-stationary dynamics make deterministic graph pruning unreliable.
@@
The variational parameters $\phi_{i}$ are learned by minimizing the Kullback–Leibler (KL) divergence between the approximate posterior $q(Z_{i};\phi_{i})$ and the true Bayesian posterior $p(Z_{i}\mid G^{\mathrm{env}}_{\mathcal{V}_{i}},D_{i})$:
This approximation enables each agent to learn a structured distribution over interaction graphs that reflects the most informative subgraphs for policy optimization. Following the standard variational inference framework, minimizing this KL divergence is equivalent (up to an additive constant) to maximizing the evidence lower bound (ELBO), or equivalently, minimizing the negative ELBO:
@@
where the likelihood term $\log p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})$ is instantiated by the graph-conditioned actor loss $\mathcal{L}_{\theta,\varphi}$ defined in Definition 4, which provides task feedback to guide the learning of latent structures. (See Appendix A.3 for details.)
@@
$\mathcal{L}_{\mathrm{ELBO}}=\mathbb{E}_{q(Z_{i};\phi_{i})}\!\left[-\mathcal{L}_{\theta,\varphi}\right]-\sum_{j\in\mathcal{N}_{i}}\mathrm{KL}\!\left(q(Z_{ij};\phi_{ij})\|p(Z_{ij})\right),$ (11)
where $\mathcal{L}_{\theta,\varphi}$ denotes the graph-conditioned policy loss defined in Definition 4, $q(Z_{ij};\phi_{ij})=\mathrm{Bernoulli}(\sigma(\phi_{ij}))$ is the variational edge distribution, and $p(Z_{ij})=\mathrm{Bernoulli}(\lambda)$ is the Bernoulli prior with retention bias $\lambda$. This objective jointly optimizes the policy and the structure of the latent subgraph via variational inference.
@@ 5.1 Experimental Setup
We evaluate BayesG on five benchmark scenarios for adaptive traffic signal control (ATSC), implemented using the SUMO microscopic traffic simulator [38]. Each environment simulates peak-hour traffic, with one MDP step corresponding to a fixed control interval. Agents observe traffic conditions via induction-loop detectors (ILDs), including vehicle density, queue lengths, and waiting times on incoming lanes, and control local traffic signals. The reward is the negative number of halted vehicles, normalized by a fixed scale. These environments naturally satisfy Definition 1: state transitions depend on local and neighbor actions, rewards are localized per agent, and neighborhood structure is fixed by the road network. Detailed MDP mappings are in Appendix B.  [^1] Code and data are available at https://github.com/Wei9711/BayesG.
@@
NewYork. To assess scalability, we introduce three large-scale scenarios—NewYork33, NewYork51, and NewYork167—comprising 33, 51, and 167 signalized intersections, respectively. These networks are derived from real-world Manhattan layouts (see Appendix C for more details). We use a 20-second control interval (increased to 40 seconds for NewYork167) and simulate 500 steps per episode. Intersections feature heterogeneous phase designs and ILDs configurations. States include normalized lane-level metrics and neighborhood-aware features. These environments pose significant challenges for coordination and generalization in decentralized MARL.
@@ 5.1.1 Baselines
ConseNet [16]: A consensus-based method where critics synchronize neighbours’ parameters via local averaging to promote stability and coordination, without exchanging message content.
@@ 5.4 Ablation Studies
Mask Input Features. We further examine how different information types affect the learned mask. Specifically, we use:(1) State: Mask is generated based on neighbor observations $s_{\mathcal{V}_{i}}$. (2) Trajectory: Uses LSTM hidden states $h_{\mathcal{V}_{i}}$, capturing temporal behavior. (3) Policy: Based on neighbor action distributions $\pi_{\mathcal{N}_{i}}$. Results in Figure 5 (right) indicate that trajectory-based features lead to stronger performance that state-only, as the incorporate richer temporal context. Policy-based masks also provide meaningful signals, reflecting agents’ behavioral intentions. The best results are achieved when the learned mask uses a combination of these features, as implemented in BayesG.
@@ 6.1 Comparison with Sampling-based Networked MARL
Recent work has explored sampling strategies for scalability in networked MARL, but differs fundamentally in what is sampled, how distributions are determined, and where adaptivity occurs.
Sampling targets and distributions. Qu et al. [26] establish convergence results for networked systems with local dependencies on fixed graphs. Lin et al. [23] sample active links from a predefined distribution $\mathcal{D}$ for $\mu$-decay analysis. Anand and Qu [27] use uniform agent subsampling ($k$ out of $n$) for a global controller. In contrast, our method learns context-dependent, per-edge distributions $q(Z_{i};\phi_{i})$ optimized jointly with the policy via an ELBO.
Locus of adaptivity. Lin et al. [23] achieve dynamics-level stochasticity through time-varying link sets. Anand and Qu [27] vary agent subsets for a global controller. Our approach operates at the policy level: each agent samples a task-adaptive 1-hop subgraph for local decision-making.
First-hop inclusion probabilities. Lin et al. [23] use the marginal of $\mathcal{D}$ (predefined, not state-adaptive). Anand and Qu [27] use uniform inclusion ($k/n$). Our method learns edge-specific probabilities $\Pr[z_{ij}=1]=\sigma(\phi_{ij}(\cdot))$ that are state and trajectory-dependent.
These approaches provide valuable convergence guarantees under structural assumptions. Our contribution is complementary: learning which neighbors to attend to, optimized end-to-end with the policy. A detailed comparison is in Table 2 (Appendix H).
@@ 6.2 Distinction from Dec-POMDP-based Coordination Graphs
In contrast, Dec-POMDP-based methods such as CASEC [15], DCG [33], DICG [29], SOP-CG [34], and GACG [22] assume global state $s$, global reward $R(s,\mathbf{u})$, joint transitions $p(s^{\prime}|s,\mathbf{u})$, and centralized training with decentralized execution (CTDE). They allow unrestricted graph rewiring unconstrained by physical topology, making them incompatible with physically grounded networked systems. Our baselines—CommNet [40], NeurComm [7], and LToS [23]—communicate only with physical neighbors, respect fixed topology, and perform fully decentralized learning. While Dec-POMDP coordination graphs are important contributions, we exclude them due to incompatibility with Spatiotemporal-MDP’s decentralized, physically constrained nature.
@@ 7 Conclusion
We presented BayesG, a variational framework for learning latent interaction graphs in networked multi-agent reinforcement learning. By operating over ego-graphs—localized subgraphs centered on each agent—BayesG enables decentralized agents to infer task-adaptive communication structures using local observations and Bayesian inference. Our approach integrates graph inference with actor–critic training via an ELBO objective, jointly optimizing both the interaction topology and policy. Experiments on five adaptive traffic signal control benchmarks demonstrate that BayesG consistently outperforms strong baselines, achieving superior coordination and reduced congestion. Case studies and ablation results further highlight the efficiency and interpretability of the learned graphs. BayesG offers a principled approach to structure-aware cooperation in decentralized settings and opens up new opportunities for scalable learning in networked systems.
@@ A.3 Derivation of Variational Objective and ELBO for BayesG
$=\mathbb{E}_{q(Z_{i})}\left[\log q(Z_{i})-\log p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})-\log p(Z_{i})\right]+\log p(D_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}})$ (26)
$=-\mathbb{E}_{q(Z_{i})}\left[\log p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})+\log p(Z_{i})-\log q(Z_{i})\right]+\log p(D_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}})$ (27)
$=-\mathbb{E}_{q(Z_{i})}\left[\log p(D_{i}\mid Z_{i},G^{\text{env}}_{\mathcal{V}_{i}})+\log p(Z_{i})-\log q(Z_{i})\right]+\text{const},$ (28)
where we have used the fact that $\log p(D_{i}\mid G^{\text{env}}_{\mathcal{V}_{i}})$ does not depend on $Z_{i}$ or $\phi_{i}$. Therefore, minimizing the KL divergence is equivalent (up to an additive constant) to maximizing the evidence lower bound (ELBO):